OpenStudy (riley):

Find the derivative of the function. s(x) = ln(4x − 3)^1.4

OpenStudy (riley):

qed your my hero >_< like srlsy you have no idea . . .

OpenStudy (anonymous):

s '(x) = (1.4(ln(4x-3))^0.4)(4/(4x-3))

OpenStudy (anonymous):

\[\frac{10 \log ^{\frac{3}{2}}(4 x-3)}{4 x-3} \]